5/36
Probability of flipping a coin twice and getting tails both times?
P(tail) is 0.5. You want P(tail) AND P(tail) = 0.5 * 0.5 = 0.25 or 1/4.
What is a year with three digits?
Any year between and including A.D. 100 and 999. So, A.D. 387, 234, etc. Or, any year between B.C. 100 and 999.!
What is an example of the address of a set of data in a spreadsheet?
For example, to identify a rectangle that begins with cell A1 and extends down to cell D3, one address for the range would be A1:D3.
How many addition sgins should be put between digits of the number 987654321?
It depends on what you want to do.
How many three digit numbers can you make using the one two and three?
it would be
321
312
231
213
123
132
How did pascal make pascal's triangle?
pascal drew a triangle, and then signed his name at the bottom of the page. He was 7 years old
I suppose you mean "proton number". That refers to the number of protons in the nucleus of an atom.
3
How many 3 digits numbers can be formed from numbers 3-7?
you can choose from 3, 4, 5, 6 and 7 and can place them on 3 digits
so 5 options over 3 digits. so 5*5*5 = 125 numbers
=================================
If you're not allowed to use the same digit twice, then
The first digit can be any one of the five. For each choice ...
The second digit can be any one of the remaining four. For each choice ...
The third digit can be any one of the remaining three.
Total number of different arrangements = 5 x 4 x 3 = 60 .
How many 2 digit numbers have no 2 at all?
72. (with the range of two digit numbers being from 10 to 99).
What fields you can take after passing ssc in general maths?
We can any field except engineering....all d best
is defined as the condition of being without children. Childlessness may have personal, social and/or political significance. There are two distinguishable types of childlessness, voluntary and involuntary. Voluntary childlessness, also described as childfree, is a consequence of having made a decision not to reproduce. To be childless not by choice is defined as involuntary childlessness.
666
If the mean score on an exam is 88 and the z-score for 74 is -1.5 what is the standard deviation?
Z = (X - M)/SD
So SD = (X - M)/Z = (74 - 88)/(-1.5) = 14/1.5 = 9.333....